Two dice are rolled. Find the odds against a sum of 5.
Answer by stanbon(57307)

Two dice are rolled. Find the odds against a sum of 5.
-----
Ways to get sum of 5: 1/4,2/3,3/2,4/1-----4 ways
----
P(sum of 5) = 4/36 = 1/9
P(not sum of 5) = 8/9
----
Odds against E = P(not E)/P(E)
-----
Probability against sum of 5 = (8/9)/(1/9) = 8:1
========================================================
Cheers,
Stan H.
===================